OpenStudy (gabylovesyou):

What is the value of x in the solution to the following system of equations? x − 2y = 2 3x + y = 6

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (anonymous):

you could use substitution!! :) 1. if 3x+y=6 2. y = 6-3x you can substitute this 6-3x for the y value in x − 2y = 2 3. x - 2(6-3x) = 2 4. x - 12 + 6x = 2 5. 7x = 14 6. x = 2 the value of x is 2! hope that helps!!!

OpenStudy (gabylovesyou):

thank you !!! :D

OpenStudy (anonymous):

honestly anytime!!!! :)